OpenStudy (anonymous):

find the general solution of sin 2x+cos x = 1

OpenStudy (anonymous):

sin2x = 1-cosx 2sinxcosx = 2(sin(x/2))^2 4sin(x/2) cos(x/2)cosx - 2(sin(x/2))^2 =0 2sin(x/2) ( 2cos(x/2)cosx - sin(x/2) ) =0 we may get one solution from 2sin(x/2) =0 or x/2 = n pi or x=2n pi and further...

OpenStudy (anonymous):

one solution is given as (2n+1)pi/2 and another is n(pi) + (-1)^n 7pi / 6 how do i get these ? @matricked

OpenStudy (anonymous):

is your question with sin2x or (sin x)^2

OpenStudy (anonymous):

is is sin 2 x as in sin (two x )

OpenStudy (anonymous):

i think it means (sin x)^2 well if it is so then we can rewrite the question as (sin x)^2 +cosx =1 or 1-(cos x)^2x +cosx =1

OpenStudy (anonymous):

it is not sin ^2 x but sin (2x) + cos x = 1

OpenStudy (anonymous):

by the way if we put n=0 in the answer provided by you the solutions are two of the solutions pi/2 and pi/6 but none of them satisfies sin (2x) + cos x = 1

OpenStudy (anonymous):

i meant 7 pi/6

OpenStudy (anonymous):

can we proceed like this 2 sin x cos x + cos x = 1 or cos x (2 sin x+1)= 1 hence either cos x =1 or 2sin x +1 = 1 ?

OpenStudy (anonymous):

no you can't

OpenStudy (anonymous):

as a* 1/a = 1 for all a<>0